OpenStudy (anonymous):

The perimeter of some traffic sign must be 120 inches. Also, its length is 8 inches longer than its width. Find the dimensions of this sign???? Help?

OpenStudy (anonymous):

width be x lenght =8 inches longer than its width= x+8 perimeter is 120=2*(width+length) 2*(x+x+8)=120 (2x+8)=60 2x+8=60 2x=60-8 2x=52 x=26 inches widht=26inches length=width+8=26+8=34 inches

OpenStudy (anonymous):

awesome....thanks